OpenStudy (megatron):

Explain how to convert standard form to vertex form? 1) “To pass by me you must tell me how to convert standard form into the general, vertex form... I have a test on it next week.” Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. f(x) = x^2 + 6x – 16

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (megatron):

I think so. When you multiply a binomial by itself, right?

OpenStudy (jdoe0001):

well, yes, but a perfect square trinomial has.... 3 terms that look like this \(\bf a^2\pm {\color{red}{ 2}}ab+b^2\) <-- notice the the middle term the middle term is composed of the root of the two extremes times 2

OpenStudy (jdoe0001):

so.. it ends up at -> \(\bf a^2\pm {\color{red}{ 2}}ab+b^2\implies (a\pm b)^2\)

OpenStudy (megatron):

Ok so I need to rearrange f(x) into (a plus or minus b)^2?

OpenStudy (jdoe0001):

so let's take a peek at the equation, and lets firstly do some grouping \(\bf f(x) = x^2 + 6x - 16\implies f(x) = (x^2 + 6x) - 16 \\ \quad \\ \implies f(x) = (x^2 + 6x+{\color{red}{ \square }}^2) - 16\) so what do you think we need there to make that a perfect square trinomial?

OpenStudy (megatron):

6x^2?

OpenStudy (jdoe0001):

well.... notice the middle term, keep in mind the middle term is using a factor of "2" if you remove that factor of 2, what would you be left?

OpenStudy (megatron):

3?

OpenStudy (jdoe0001):

yeap, 3 thus 2*3*x = 6x keep in mind that, all we're doing is borrowing from "0", zero so if we ADD \(3^2\) we also have to SUBTRACT \(3^2\) thus \(\bf f(x) = x^2 + 6x - 16\implies f(x) = (x^2 + 6x) - 16 \\ \quad \\ f(x) = (x^2 + 6x+{\color{red}{ 3 }}^2) - 16-{\color{red}{ 3 }}^2\implies f(x)= (x+3)^2-16-9 \\ \quad \\ \implies f(x)= (x+3)^2-25\)

OpenStudy (megatron):

So how would I put this in complete sentences? f(x)=x^2+6x-16 can be converted to vertex form by grouping it to make a perfect square trinomial to get f(x)=(x^2+6x+3^2)-16-3^2 by adding 3^2 and also subtracting 3^2 to get f(x)=(x+3)^2-25 ?

OpenStudy (jdoe0001):

yeap

OpenStudy (megatron):

Thank you!

OpenStudy (jdoe0001):

yw